# Phase error rate estimation in QKD with imperfect detectors

### v2

@@ F.0.1 Bounding the smooth min entropy
$\leq\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}\geq\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}$
$+\frac{1}{2}2^{-\frac{1}{2}\left(n_{K,{1}}\left(1-h\left(\mathcal{B}_{e}\left(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\right)\right)\right)-l(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}-\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})-\log(2/\varepsilon_{\text{EV}}))\right)}\bigg{)}$
$=\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}\geq\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}+\varepsilon_{\text{PA}}\bigg{)}$
$\leq\varepsilon_{\text{PA}}+2\sqrt{\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\\ n_{K,{1}}\geq\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\kappa(\widetilde{\Omega})}$
$=\varepsilon_{\text{PA}}+2\sqrt{\Pr(\bm{e^{\text{key}}_{X,{1}}}\geq\mathcal{B}_{e}(\bm{e^{\text{obs}}_{X,\mu_{\vec{k}}}},\bm{n_{X,\mu_{\vec{k}}}},\bm{n_{K,\mu_{\vec{k}}}})\quad\land\quad\bm{n_{K,{1}}}\geq\mathcal{B}_{1}(\bm{n_{K,\mu_{\vec{k}}}}))}.$
Here, we used the leftover-hashing lemma [2, Proposition 9] on the sub-normalized state $(\rho_{|\widetilde{\Omega}})_{\wedge\Omega_{\text{EV}}}$ for the first inequality, and [2, Lemma 10] to get rid of the sub-normalized conditioning ($\wedge\Omega_{\text{EV}}$) in the smooth min entropy term in the second inequality. We used [54, Lemma 6.8] to split off the error-correction information ($\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$) and error-verification information ($\log(2/\varepsilon_{\text{EV}})$) in the third inequality. We used the bound on the smooth min entropy from Eq. 105 for the fourth inequality. The fifth equality follows by replacing the values of $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ and $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ from Eq. 53. We use the concavity of the square root function and Jensen’s inequality for the sixth inequality to pull the sum over the events and the probability inside the square root, while the seventh equality follows simply from the definition of conditional probabilities and $\kappa$ (Eq. 103).
Combining our bounds on $\Delta_{1}$ and $\Delta_{2}$ from Eqs. 109 and 108, we obtain
@@ G.2 Computing $\delta_{1}$
$\norm{\sqrt{{F}_{(b),(\text{con})}}G^{\text{con}}_{(X),(\neq)}}_{\infty}=$ $\norm{\sqrt{{F}_{(b),(\text{con})}}}_{\infty}\norm{G^{\text{con}}_{(X),(\neq)}}_{\infty}\leq 1.$ (118)
@@
where Eq. 120 is a consequence of the triangle inequality. Let us focus on a fixed value of $N=N_{0}+N_{1}\neq 0$. In this case, combining Eqs. 117, 118 and 120 allows us to obtain
@@
where the final inequality above follows from the monotonicity of the expression inside the modulus with respect to $\eta_{b_{0}},\eta_{b_{1}},d_{b_{0}},d_{b_{1}}$. However, as described in Ref. [52, Section III C] we can renormalize detection efficiencies and treat the common loss in detectors as part of the channel. Thus, without loss of generality, we consider one of the detectors to be lossless. This can be thought of as setting $\eta_{\text{max}}\rightarrow 1$ and $\eta_{\text{min}}\rightarrow\eta_{\text{min}}/\eta_{\text{max}}$. This significantly simplifies the calculations. Combining Eqs. 116, 118, 120, 121 and 114, we obtain
@@ H.1 Computing $\delta_{1}$
Similar to the analysis performed in Section G.2, we reduce the problem to bounding $\norm{\sqrt{{F}_{(Z),(\text{con})}}-\sqrt{{F}_{(X),(\text{con})}}}_{\infty}$ through Eqs. 117 and 118. Once again we exploit the block-diagonal structure in the POVM elements to compute the infinity norm $\delta_{1}^{(N)}$ of each block with total photon-number $N$ separately as done in Eq. 114. First, note that $\delta_{1}^{(0)}=\delta_{1}^{(1)}=0$. To compute $\delta^{(N)}$ for $N\geq 2$, we define
@@
for all $N\geq 2$. Once again we use the argument in [52, Section III C] to treat the common loss in detectors as part of the channel. This is equivalent to setting $\eta_{\text{max}}\rightarrow 1$ and $\eta_{\text{min}}\rightarrow r_{\eta}=\eta_{\text{min}}/\eta_{\text{max}}$. (Actually this sets $\eta_{\text{mult}}\rightarrow 1$ and one of $\eta_{i}$ to $1$ and the other to $r_{\eta}$). Thus we obtain

### v3

@@ F.0.1 Bounding the smooth min entropy
$\leq\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}>\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}$
$+\frac{1}{2}2^{-\frac{1}{2}\left(n_{K,{1}}\left(1-h\left(\mathcal{B}_{e}\left(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\right)\right)\right)-l(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})-\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})-\log(2/\varepsilon_{\text{EV}}))\right)}\bigg{)}$
$=\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}>\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}+\varepsilon_{\text{PA}}\bigg{)}$
$\leq\varepsilon_{\text{PA}}+2\sqrt{\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\\ n_{K,{1}}>\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\kappa(\widetilde{\Omega})}$
$=\varepsilon_{\text{PA}}+2\sqrt{\Pr(\bm{e^{\text{key}}_{X,{1}}}\geq\mathcal{B}_{e}(\bm{e^{\text{obs}}_{X,\mu_{\vec{k}}}},\bm{n_{X,\mu_{\vec{k}}}},\bm{n_{K,\mu_{\vec{k}}}})\quad\land\quad\bm{n_{K,{1}}}>\mathcal{B}_{1}(\bm{n_{K,\mu_{\vec{k}}}}))}.$
Here, we used the leftover-hashing lemma [2, Proposition 9] on the sub-normalized state $(\rho_{|\widetilde{\Omega}})_{\wedge\Omega_{\text{EV}}}$ for the first inequality, and [2, Lemma 10] to get rid of the sub-normalized conditioning ($\wedge\Omega_{\text{EV}}$) in the smooth min entropy term in the second inequality. We used [65, Lemma 6.8] to split off the error-correction information ($\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$) and error-verification information ($\log(2/\varepsilon_{\text{EV}})$) in the third inequality. We used the bound on the smooth min entropy from Eq. 106 for the fourth inequality. The fifth equality follows by replacing the values of $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ and $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ from Eq. 53. We use the concavity of the square root function and Jensen’s inequality for the sixth inequality to pull the sum over the events and the probability inside the square root, while the seventh equality follows simply from the definition of conditional probabilities and $\kappa$ (Eq. 104).
Combining our bounds on $\Delta_{1}$ and $\Delta_{2}$ from Eqs. 110 and 109, we obtain
@@ G.2 Computing $\delta_{1}$
$\norm{\sqrt{{F}_{(b),(\text{con})}}G^{\text{con}}_{(X),(\neq)}}_{\infty}\leq$ $\norm{\sqrt{{F}_{(b),(\text{con})}}}_{\infty}\norm{G^{\text{con}}_{(X),(\neq)}}_{\infty}\leq 1.$ (119)
@@
where Eq. 121 is a consequence of the triangle inequality. Let us focus on a fixed value of $N=N_{0}+N_{1}\neq 0$. In this case, combining Eqs. 118, 119 and 121 allows us to obtain
@@
where the final inequality above follows from the monotonicity of the expression inside the modulus with respect to $\eta_{b_{0}},\eta_{b_{1}},d_{b_{0}},d_{b_{1}}$. However, as described in Ref. [59, Section III C] we can renormalize detection efficiencies and treat the common loss in detectors as part of the channel. Thus, without loss of generality, we consider one of the detectors to be lossless. This can be thought of as setting $\eta_{\text{max}}\rightarrow 1$ and $\eta_{\text{min}}\rightarrow\eta_{\text{min}}/\eta_{\text{max}}$. This significantly simplifies the calculations. Combining Eqs. 117, 119, 121, 122 and 115, we obtain
@@ H.1 Computing $\delta_{1}$
Similar to the analysis performed in Section G.2, we reduce the problem to bounding $\norm{\sqrt{{F}_{(Z),(\text{con})}}-\sqrt{{F}_{(X),(\text{con})}}}_{\infty}$ through Eqs. 118 and 119. Once again we exploit the block-diagonal structure in the POVM elements to compute the infinity norm $\delta_{1}^{(N)}$ of each block with total photon-number $N$ separately as done in Eq. 115. First, note that $\delta_{1}^{(0)}=\delta_{1}^{(1)}=0$. To compute $\delta^{(N)}$ for $N\geq 2$, we define
@@
for all $N\geq 2$. Once again we use the argument in [59, Section III C] to treat the common loss in detectors as part of the channel. This is equivalent to setting $\eta_{\text{max}}\rightarrow 1$ and $\eta_{\text{min}}\rightarrow r_{\eta}=\eta_{\text{min}}/\eta_{\text{max}}$. (Actually this sets $\eta_{\text{mult}}\rightarrow 1$ and one of $\eta_{i}$ to $1$ and the other to $r_{\eta}$). Thus we obtain